### v3

Abstract
Eavesdropping has been a long-standing threat to the security and privacy of wireless communications, since it is difficult to detect and costly to prevent. As networks evolve towards Sixth Generation (6G) and semantic communication becomes increasingly central to next-generation wireless systems, securing semantic information transmission emerges as a critical challenge. While classical physical layer security (PLS) focuses on passive security, the recently proposed concept of physical layer deception (PLD) offers a semantic encryption measure to actively deceive eavesdroppers. Yet the existing studies of PLD have been dominantly information-theoretical and link-level oriented, lacking considerations of system-level design and practical implementation.
In this work we propose Visual ENcryption for Eavesdropping NegAtion (VENENA), an artificial intelligence-enabled framework for secure image-based communication. VENENA protects confidential messages by encoding them visually while actively deceiving eavesdroppers: legitimate receivers use artificial intelligence (AI)-based classifiers to extract true message semantics, while interceptors perceive only falsified content. The framework transmits two superimposed image components with different power levels–a high-power decoy image and a low-power correction mask–ensuring only authorized receivers with favorable channel conditions can reconstruct the true message. Experimental validation demonstrates over 93% accuracy for legitimate users while limiting eavesdropper success to 52% even when system design is fully known, validating VENENA’s active defense capability for 6G semantic communication.
{IEEEkeywords}
Cyber security, physical layer deception, visual encryption, image poisoning
1 INTRODUCTION
While the rapid development and evolution of wireless technologies have contributed to the latest Fifth Generation (5G) mobile systems, and leading us towards the future Sixth Generation (6G) networks, the imperative for robust security and privacy measures has also become increasingly pronouced. As outlined in [1], the rapid pace of technological advancements has complicated the cybersecurity landscape, necessitating transformative security strategies to counteract both existing and emerging threats effectively.
One of the major concerns in wireless security and privacy is that the classical cryptographic algorithms are no more guaranteed to be secure in the future. On the one hand, the recent advent of quantum computing poses a significant threat to such methods of computational security. On the other hand, the extensive softwarization, virtualization, and cloudification of network functions in 5G and future 6G networks introduce new vulnerabilities in multi-tenant environments. These increase the risk of data leakage, further degrading the security provided by classical cryptographic methods that fundamentally rely on confidential authentication information. In this context, the concept of physical layer security (PLS) has emerged as a promising alternative to traditional cryptographic methods. Leveraging the physical properties of the wireless channel to provide secure communication, PLS offers a new paradigm from preventing eavesdropping attacks and ensuring data confidentiality.
However, despite enhancing passive security, conventional PLS technologies cannot actively defend against eavesdroppers–they lack the capability to discourage, detect, or disrupt adversaries. In this regard, PLS performs no better than classical cryptographic security measures. To bridge this gap, the concept of physical layer deception (PLD) was first proposed in [2], which introduced deceptive signaling via non-orthogonal multiplexing to mislead eavesdroppers under finite blocklength constraints. This idea was later extended in [3] to derive secrecy capacity bounds and optimal power-allocation strategies, and further applied to orthogonal frequency-division multiplexing (OFDM) systems in [4]. By deceiving eavesdroppers with falsified information, while simultaneously transmitting the true information to legitimate receivers, PLD provides an active defense mechanism to enhance the security of wireless communications.
Especially, the practical value of PLD emerges in scenarios where adversaries can suffer from inaccurate information. For example, in military tactical communications, forward units can transmit target coordinates as deceptively-encoded messages: intercepting adversaries decode falsified coordinates that misdirect their resources, while legitimate receivers reconstruct true coordinates. In confidential business communications, as another instance, corporate executives can transmit merger target identifiers or product launch schedules in the same way: malicious insiders or competitors with leaked protocols intercept falsified information that triggers counterproductive actions, while authorized partners receive true intelligence. Such active deception proves more damaging to adversaries than signal denial, as they commit resources based on false intelligence rather than merely lacking information. We have later proven that PLD is a semantic encryption solution for secure transmission, which relies on the secrecy of the semantic knowledge instead of that of the messages [5].
As the core part of the PLD framework, the deceptive ciphering algorithm is required to fulfill two specific requirements. First, the codebooks of ciphertext and plaintext have to share a same codewrd set, i.e. every possible ciphertext codeword must also be a valid plaintext codeword. Second, every individual codeword must contain meaningful semantic information. While the preliminary works on PLD have been only focusing on the link-level design, little has been discussed about the application scenario or the system-level design, which can be constrained by the above requirements. essential semantic encoding mechanisms to generate meaningful semantic codewords. This creates a critical implementation gap between theoretical PLD concepts and practical semantic communication systems.
Visual encryption provides the essential semantic encoding capability that bridges this gap between theoretical PLD requirements and practical implementation. By mapping confidential messages onto semantically meaningful visual content, visual encryption enables the creation of dual-purpose codewords that satisfy both PLD’s semantic richness requirement and the deception objectives of active security. However, traditional visual encryption approaches remain fundamentally passive, lacking the active deception capabilities needed to counter sophisticated eavesdroppers who may possess full system knowledge.
As a proof of concept for PLD, in this work we propose a deceptive visual encryption framework called Visual ENcryption for Eavesdropping NegAtion (VENENA), which represents the first system-level implementation of PLD theory by combining PLD with image poisoning to realize secure and deceptive transmission of secrecy messages over wireless wiretap channels. VENENA advances the field by:
  1. 1.
    System-level instantiation: While prior PLD work [3, 5] established information-theoretic foundations deriving secrecy capacity bounds under the assumption that de- ceptive ciphering algorithms exist, no concrete implementation of such algorithms or system architecture was provided. VENENA provides the first practical realization of PLD by addressing: (i) semantic encoding via visual classification networks, (ii) deceptive content generation through gradient matching attacks, and (iii) power-domain multiplexing for dual-message transmission, transitioning PLD from theoretical abstraction to engineered system;
  2. 2.
    Artificial intelligence (AI)-enabled deceptive content generation: We introduce gradient matching attack (GMA)-based image poisoning as the realization of PLD’s deceptive ciphering requirement, enabling adversarial perturbations that maintain visual similarity while achieving semantic transformation. This addresses the critical gap between PLD’s requirement for “semantically meaningful codewords” and practical generation mechanisms;
  3. 3.
    Empirical validation against informed adversaries: We demonstrate quantifiable security gains ($40\%$ reduction in eavesdropper accuracy) even when adversaries possess full system knowledge, validating PLD’s active defense capability against insider threats that defeat traditional visual encryption.
An overview of VENENA’s features compared to conventional PLS and existing PLD works is provided in Table 1.
Table 1: Overview of VENENA features compared to conventional PLS and existing PLD works.
Conventional PLS Existing PLD VENENA
Active defence No Yes Yes
Basis Randomness of the physical channel Statistical channel superiority of the legitimate receiver over eavesdropper Statistical channel superiority of the legitimate receiver over eavesdropper
Semantic encryption Irrelevant Neccesary, unspecified AI-driven visual encryption
System-level demonstration Reported Unavailable so far Yes (first practical PLD framework)
The remainder of this article is organized as follows. In Sec. 2 we provide a brief tutorial to the related works, outlining the state-of-the-art in the fields of PLS, cyber deception, PLD, visual encryption, and image poisoning. In Sec. 3 we introduce the proposed VENENA framework, detailing its principles and design. As the technical enabler, the image poisoning algorithm is presented in Sec. 4, with a step-by-step description of the poisoning process. Subsequently, in Sec. 5 we present the numerical verification of the VENENA framework, demonstrating its effectiveness in secure and deceptive wireless image transmission. Some complementary discussions regarding the efficiency aspects of our proposed solution are made then in Sec. 6. In the end, with Sec. 7 we conclude the article and give a glance into future research possibilities.
2 RELATED WORKS
II-A PHYSICAL LAYER SECURITY
Research on PLS can be traced back to Wyner’s foundational research [6], which expanded Shannon’s perfect secrecy theorem [7] by introducing quantifiable strong secrecy metrics for wiretap channel scenarios. This breakthrough led to extensive studies examining secrecy characteristics across diverse wiretap channel configurations: symmetric and degraded channels, fading systems, multi-antenna and relay architectures, and broadcast/multiple-access scenarios. Building upon this theoretical foundation, comprehensive surveys have systematically classified PLS techniques across different wireless technologies. The work in [8] provides a complete taxonomy of physical layer security approaches for confidentiality, categorizing techniques from information-theoretic perspectives to practical implementations in modern wireless systems. Similarly, [9] offers an extensive analysis of optimization approaches specifically targeting wireless physical layer security, encompassing resource allocation strategies, precoding techniques, and cooperative protocols. The theoretical understanding gained from these investigations has enabled various security-enhancing techniques: resource allocation in radio networks, precoding and beamforming strategies, antenna and node selection, cooperative protocols, and specialized channel coding. These techniques collectively strengthen communication system security. However, despite these advances, conventional PLS approaches remain fundamentally passive, lacking the capability to actively counter sophisticated eavesdroppers.
II-B CYBER DECEPTION
The application of deception in cybersecurity was first popularized through Mitnick’s notable social engineering exploits [10]. This concept eventually evolved into defensive mechanisms, initially emerging as honeypot systems before expanding into a comprehensive suite of deception-based security solutions. The field has matured significantly over the past decades, with a recent comprehensive survey [11] providing detailed analysis of deception techniques specifically designed for honeypots and honeynets. The survey reveals that modern cyber deception technologies have evolved from simple decoy systems to sophisticated frameworks capable of adaptive behavior and real-time response to attackers. These technologies operate on a fundamental strategy of deploying decoy targets, such as synthetic data designed to mimic confidential information, to misdirect and confuse potential attackers while safeguarding authentic assets. Additionally, these deceptive methods serve as effective tools for exposing adversaries, enabling a proactive security stance. Throughout the last thirty years, deception-based security has matured significantly, with implementations spanning across all major layers of information systems: network, system, application, and data. Researchers have developed numerous approaches to address cyber threats through detection, prevention, and mitigation strategies. The evolution from passive honeypots to active deception frameworks demonstrates the critical shift from merely detecting attacks to actively misleading and disrupting adversaries. For an in-depth exploration of current developments in deception technologies, interested readers should consult [12].
II-C SEMANTIC TRANSMISSION SECURITY
The emerging concept of semantic communication has started reshaping the understanding and designing of communication systems in various aspects. In the context of security, semantic communication can be leveraged by different means, such like semantic data encryption, covert communication, and quantum-empowered semantic security. A comprehensive survey is provided by [13].
II-D PHYSICAL LAYER DECEPTION
Deception technologies at the wireless physical layer remain an emerging field of study. The original concept of PLD was first proposed in [2], which introduced a non-orthogonal multiplexing-based approach to embed deceptive information directly into the physical layer, thereby misleading eavesdroppers under finite blocklength constraints. Subsequent work in [3] generalized this framework, establishing its theoretical foundations through secrecy-capacity analysis, while [4]further demonstrated its applicability to OFDM systems, bridging the gap between theoretical modeling and practical waveform design. More recently, [5] established PLD formalized PLD as a semantic encryption mechanism, highlighting its potential for secure semantic communication in next-generation wireless networks. In a closely related but independent line of research on waveform-level deception, similar ideas were explored in  [14], where researchers exploit multi-input multi-output (MIMO) systems’ spatial diversity to generate a secondary deceptive signal stream alongside the legitimate one. However, these studies primarily focus on spatial deception rather than semantic deception, and therefore lack the systematic framework necessary for practical deployment in semantic communication scenarios.
These efforts represent early steps toward applying deception principles at the physical layer, but they do not address the semantic encoding requirements essential for meaningful and controllable deception in future 6G systems.
Despite these advances, existing PLD research remains predominantly information-theoretic, concentrating on link-level secrecy analysis under the assumption that deceptive ciphering algorithms exist with desired properties. No prior work has demonstrated a concrete architecture or implementation capable of generating semantically meaningful falsified content. The VENENA framework fills this critical gap by providing the first system-level realization of PLD through AI-enabled image poisoning, power-domain multiplexing, and empirical performance validation. In contrast to the abstract and asymptotic analyses in previous PLD studies, VENENA delivers a complete end-to-end system architecture, quantifiable performance results, and experimental verification under realistic channel conditions.
II-E VISUAL ENCRYPTION
Visual encryption is a technique that aims to protect the confidentiality and privacy of visual information by embedding it into images in a visually unrecognizable form. The goal is to ensure that the encrypted image appears similar to the original image while concealing the embedded information. Various approaches have been proposed in the literature to achieve visual encryption through information embedding.
Digital watermarking is one of the most well-known techniques for embedding information into images [16]. However, the amount of information that can be embedded into an image using digital watermarking is limited, and increasing the embedding capacity often comes at the cost of reduced robustness or visual quality.
Optical transform-based approaches have also been explored to achieve visual encryption through mathematical transformations. For example, quaternion Fresnel transform combined with chaotic systems can process multiple images holistically and transform recognizable visual content into noise-like patterns suitable for watermarking and secure transmission [17]. While such methods provide strong obfuscation through mathematical operations, they fundamentally rely on key secrecy–when adversaries obtain system parameters through protocol leakage or insider access, the encrypted content appears merely as random noise without providing active misdirection.
Recent advancements in visual encryption have leveraged deep learning techniques to embed information into images. Deep learning-based approaches utilize the power of convolutional neural networks to learn complex embedding and extraction mappings, enabling more robust and adaptive visual encryption schemes [18]. However, these systems remained primarily defensive and lacked deceptive capabilities needed in adversarial environments. The critical breakthrough came with the integration of adversarial machine learning techniques, particularly image poisoning, which enables dual functionality: securely transmitting hidden information while actively misleading potential adversaries through carefully crafted perturbations. Image poisoning with embedded information, it becomes possible to transmit secret messages or watermarks securely while also deceiving potential adversaries [19]. Unlike purely defensive approaches, poisoning techniques enable dual functionality: securely transmitting hidden information while actively misleading potential adversaries through carefully crafted perturbations. This advancement overcame the passive nature of previous methods, creating systems capable of both concealment and misdirection.
3 THE VENENA FRAMEWORK
We consider an wiretap channel model with a sender (Alice), a legitimate receiver (Bob), and an eavesdropper (Eve). Alice wants to transmit a short confidential message (e.g., a number) to Bob, while preventing Eve from correctly recovering the message.
(a) Naive approach: Alice maps messages to semantic tags and transmits corresponding images. Bob uses a classification network to decode. Security relies entirely on keeping the semantic mapping table secret from Eve, failing when Eve is an insider or protocols are standardized.
(b) VENENA: Alice generates a falsified message encoded in an original image (deceptive content), then creates a poisoned version that classifies as the true message. The poison mask (difference between images) and original image are power-multiplexed and transmitted. Bob decodes both components with favorable channel conditions and reconstructs the poisoned image to recover the true message. Eve decodes only the high-power original image due to worse channel conditions, perceiving the falsified message. VENENA remains secure even with full system knowledge, as the poison mask provides semantic confidentiality.
Figure 1: Visual encryption of confidential messages: 1(a) naive design to defend eavesdroppers without essential knowledge for decryption, and 1(b) the VENENA framework to deceive insider eavesdroppers.
III-A VISUAL ENCRYPTION
To realize such confidential transmission, Alice employs a semantic codec that operates through categorical labels called semantic tags (e.g., “airplane”, “automobile”, or “bird”). Each raw message is mapped onto a semantic tag, to which an image dataset is associated. The encoding process randomly selects an image from the corresponding dataset, transforming the raw message (number) into a visual message (image). Bob then uses a pre-trained classification network to extract the semantic tag and recover the message. As long as the mapping between the numbers and the semantic tags is unknown to Eve, the semantic information remains confidential even if Eve can correctly decode the image. This visual encryption approach is illustrated in Fig. 1(a).
Obviously, the security of the visual encryption scheme completely relies on the secrecy of the mapping between the message codewords and the semantic tags. Unfortunately, there are multiple cases where the confidentiality of this information can be compromised. For example, when the protocol design is standardized and open, or when the eavesdropper is a malicious insider, or when the eavesdropper is another legitimate receiver of the system, it becomes possible to learn the mapping from statistical analysis of the eavesdropped images. In such cases, the visual encryption scheme fails to provide the desired information secrecy.
III-B DECEPTIVE VISUAL ENCRYPTION
To enhance the security of the visual encryption scheme, we propose a novel framework called VENENA, which leverages the concept of PLD and image poisoning. As shown in Fig.1(b), instead of simply sending an image with the message-associated semantic tag, Alice first randomly generates a falsified message, which is then visual-encoded into a correspondingly tagged image, which we call the original image. This image serves as a deceptive message, since it carries deceptive semantics. It is then processed with a poisoning network in such a way, that the poisoned image at output shall be classified with the true semantic tag, while remaining miminally diversed from the original image in the sense of error power. It is this poisoned image that carries the true semantics. Afterwards, a poison mask is obtained by pixel-wise differentiating the original image from the poisoned one. Seen from the semantic PLD perspective, the poison mask carries a random knowlege that maintains the confidentiality of the true semantics, when both the semantic codec and the deceptive message are leaked [5]. The original image and the poison mask are then multiplexed and sent together, supposed to be decoded by Bob separately for construction of the poisoned image. Specifically, while VENENA is agnostic to the multiplexing method, in this work we apply a power-domain non-orthogonal multiplexing, where the original image is allocated with significantly higher power than the poison mask, so that it can be first decoded treating the poison mask as a weak inteference. Subsequently, the decoded original image is subtracted from the received signal, and the poison mask can be decoded from the residual signal (known as the technique of successive interference cancellation (SIC) [15]). The power allocation shall be fine tuned in such way, that the signal-to-interference-and-noise ratio (SINR) at Bob (i) significantly exceeds the necessary level for successful decoding of the original image, but (ii) fulfills the requirement for decoding of the poison mask with only a small margin. Thus, PLS is selectively applied on the poison mask component, which carries the differential information between the true and falsified messages, while the original image carrying falsified message is well exposed to potential eavesdroppers.
With appropriate beamforming by Alice towards Bob, combined with other common measures such like secrecy protected zones, it is generally agreed that Eve in wireless eavesdropping scenarios suffers from statistically worse channel conditions than Bob does, as explained in [3]. Therefore, while the original image, due to its high power, may be well decoded by both Bob and Eve, the poison mask is only decodable by Bob. Having both components decoded, Bob will be able to regenerate the poisoned image, and therewith obtain the true message. In contrast, missing the poison mask, Eve will only perceive the falsified message carried by the original image, and therefore be deceived.
4 IMAGE POISONING
While the VENENA framework provides the overall approach for deceptive visual encryption, its effectiveness heavily relies on the image poisoning technique that enables semantic transformation while maintaining visual similarity. In this section, we present a novel targeted data poisoning attack specifically designed for this purpose.
Given a clear image $I_{A}$ belonging to the class $A$ and a target class $B$, the desired method is supposed to poison $I_{A}$ to $I_{B}^{\text{p}}$ by perturbing images a fraction of images. This poison process is accomplished by GMA, which exploits gradient descent learning: during training, each sample contributes a gradient guiding the model’s parameter updates. GMA crafts poisoned samples whose training gradients align with the adversarial gradient that would cause the target misclassification ($A\to B$). By aligning these gradient directions, the poisoner steers the model’s learning process toward the desired confusion, such that the trained model incorrectly predicts class B for the target image $I_{A}$ during inference. Critically, perturbations are constrained to remain visually imperceptible (bounded by $\epsilon$), ensuring the attack remains stealthy.
Algorithm 1 The GMA algorithm
Clean image $I_{A}$, Target class $T$, Poison budget $\epsilon$, Loss function $\mathcal{L}$, Trained model $F_{\theta}$, Hyperparameters $\lambda_{1},\lambda_{2},\lambda_{3}$
Poison $P$
Initialize $P\leftarrow C$
for $i=1,\ldots,n_{\text{iter}}$ do
  Sample a clean batch ${(x_{c},y_{c})}\sim D_{c}$
  $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(F_{\theta}(x_{c}),y_{c})$
  $g_{t}\leftarrow\nabla_{\theta}\mathcal{L}(F_{\theta}(T),y_{t})$
  $g_{p}\leftarrow\nabla_{\theta}\mathcal{L}(F_{\theta}(P),y_{p})$
  $\mathcal{L}_{\text{match}}\leftarrow-\sum{l=1}^{L}\frac{g_{t}^{l}\cdot g_{p}^{l}}{|g_{t}^{l}|2|g_{p}^{l}|2}$
  $\mathcal{L}_{\text{class}}\leftarrow\text{CrossEntropy}(F{\theta}(P),y_{t})$
  $\mathcal{L}_{\text{norm}}\leftarrow\sum{l=1}^{L}|g_{p}^{l}|2^{2}$
  $\mathcal{L}_{\text{total}}\leftarrow\lambda_{1}\mathcal{L}_{\text{match}}+\lambda_{2}\mathcal{L}_{\text{class}}+\lambda_{3}\mathcal{L}_{\text{norm}}$
  $P^{\prime}\leftarrow P-\alpha\nabla{P}\mathcal{L}_{\text{total}}$
  $P^{\prime}\leftarrow\text{clip}(P^{\prime},P-\epsilon,P+\epsilon)$
  $P\leftarrow P^{\prime}$
end for
Train the classification model $F{\theta}$ on $C\cup P$
return $P$, $F_{\theta}$
To achieve this, the poisoner computes two sets of gradients: the target gradient and the poison gradient. The target gradient represents the direction of change in the model’s parameters that would increase the loss for the target image with respect to the adversarial target label. The poison gradient, on the other hand, is the sum of the gradients of the poisoned examples with their corresponding labels.
The poisoner then minimizes the negative cosine similarity between the target gradient and the poison gradient. By aligning these gradients in the same direction, the poisoner ensures that the model’s updates during training will move in a direction that increases the likelihood of misclassifying the target image.
By iteratively updating the perturbations to minimize the negative cosine similarity between the target and poison gradients, the poisoner crafts poisoned examples that effectively manipulate the model’s behavior, causing it to misclassify the target image during inference. Throughout this optimization process, the perturbations are constrained within $\epsilon$. A pseudo code of GMA is presented in Algorithm 1.
5 NUMERICAL VERIFICATION
V-A SIMULATION SETUP
V-A-1) Data Format
Every confidential message is a random digit in $\mathcal{M}=\{0,1,2,\dots 9\}$. An image database containing $10$ classes of photos, generated from the CIFAR-10 data, is therefore used for the visual encryption, the mapping between messages and image classes (tags) is listed in Tab. 2. Every message image is $32\times 32$ pixels in RGB, with each pixel quantized to $8$ bits per channel.
Table 2: Mapping between messages and semantic tags
Message 0 1 2 3 4
Tag airplane automobile bird cat deer
Message 5 6 7 8 9
Tag dog frog horse ship truck
V-A-2) Networks Training and Evaluation
Classifier Network
Since GMA is designed to deceive specific neural networks, we must first train the implementation of the classifier network on the receiver’s side. It employs a pre-trained vision transformer architecture, trained offline on the CIFAR-10 dataset without poisoned data. This pre-training approach ensures that during operational deployment, the system requires only inference operations with minimal computational cost. The victim network is trained using the specified hyperparameters and configuration options. The training image size is set to $224\times 224$ pixels, and the batch sizes for training and evaluation are both set to 64.
The learning rate for the victim network is set to 0.03 with a cosine decay schedule. The weight decay is set to 0, and the total number of training steps is 10 000. The learning rate warmup is performed for the first 500 steps, and the maximum gradient norm is clipped to 1.0. The victim network training utilizes mixed-precision training (FP16) with an optimization level of O2.
Table 3 presents the inference time evaluation of our classifier model. With $L\cdot d=11.17$ million parameters (memory size $42.63\text{\,}\mathrm{MB}$), the model achieves a mean inference time of ${1.8866\pm 0.2258}$\mathrm{ms}$\$ per image, yielding a throughput of 530.04 images per second. The low standard deviation (12% of the mean) and stable min-max range (${1.69\--3.19}$\mathrm{ms}$\$) demonstrate consistent performance suitable for real-time applications, making it practically viable even for real-time 6G communication scenarios.
Table 3: Testbed setup and real-time performance
Metric Value
Testbed Setup
Classifier Network Parameters (#) 11 173 962
Classifier Network Size (MB) 42.63
GPU NVIDIA GeForce RTX 4090
Poisoning Time
Mean Time per Image (s) 0.3170
Standard Deviation (s) 0.0045
Min Time per Image (s) 0.3080
Min Time per Image (s) 0.3260
Throughput (images / s) 3.15
Inference Time
Mean Time per Image (ms) 1.8866
Standard Deviation (ms) 0.2258
Min Time per Image (ms) 1.6900
Min Time per Image (ms) 3.1937
Throughput (images / s) 530.04
Poisoning Process
The data poisoning process employs a gradient matching attack, which utilizes an ensemble of surrogate models that are offline trained on clean CIFAR-10 data. During the poisoning phase, these surrogate models guide an iterative optimization process that generates adversarial perturbations for 10% of samples from the source class (the penetration rate is empirically validated). The optimization aligns the gradients of poisoned samples with those that would cause misclassification of target images into the source class, while constraining perturbations within an $L_{\infty}$ bound of $\epsilon=16/255$.
Each poisoning operation requires $n_{\text{iter}}$ iterations of gradient computations across all $L$ network layers, resulting in $O(n_{\text{iter}}\cdot L\cdot d)$ complexity where $d$ represents parameter dimensionality. The layer-wise gradient matching $\mathcal{L}_{\text{match}}=-\sum_{l=1}^{L}\frac{g_{t}^{l}\cdot g_{p}^{l}}{\left|g_{t}^{l}\right|_{2}\left|g_{p}^{l}\right|_{2}}$ requires substantial memory for gradient tensor storage and manipulation.
To assess the time efficiency of the poisoning process, we carried out the poisoning on $5000$ image samples w.r.t. the classifier network trained above, with $n_{\text{iter}}=500$ iterations, and measuring the poisoning time per image. The testbed setup and test results are summarized in Table 3. As we can see, the poisoning process performs at a stationary speed with average throughput of $3.15$ images per second, which can be sufficient for low-rate transmissions of sparse confidential messages, but far from real-time for high-rate data streams.
Shall VENENA be applied in scenarios where high data rates are required, we recommend to rely on offline-poisoned image databases instead of carrying out poisoning online. With poisoned images pre-generated and stored for immediate use during transmission, reducing the time complexity from $O(n_{\text{iter}}\cdot L\cdot d)$ to $O(1)$ per image. An occoasional update of the poisoned image database shall be scheduled in this case, to eliminate long-term statistical learning by eavesdroppers.
V-A-3) Transmission Scheme
For every message $m\in\mathcal{M}$, which is one digit between 0 and 9, a distinct digit $\tilde{m}\neq m$ is randomly selected as the fake message. An image $i_{m}$ is then randomly picked from the data base $\mathcal{I}_{m}$ that corresponds to the semantic tag associated with $m$. This image is then poisoned by the network $\mathcal{P}_{m}^{\tilde{m}}$ to generate its $\tilde{m}-$induced version $i_{m}^{\tilde{m}}$. Both images are interpreted in the format of binary vectors and then their sum modulo two is calculated to obtain the binary vector of additive poison
$\left(p_{m}^{\tilde{m}}\right)_{\mathrm{b}}=\left(i_{m}\right)_{\mathrm{b}}\mathrm{XOR}\left(i_{m}^{\tilde{m}}\right)_{\mathrm{b}}.$ (1)
Both $i_{m}$ and $p_{m}^{\tilde{m}}$ are then modulated in binary phase shift keying (BPSK) into the waveform components $s_{i}$ and $s_{p}$, respectively. The two components are then mixed together with a power ratio $\alpha:(1-\alpha)$ into the transmitted signal $s$, which is transmitted at the power $P_{\Sigma}$, i.e.,
$s(t)=\sqrt{\alpha P_{\Sigma}}s_{i}(t)+\sqrt{(1-\alpha)P_{\Sigma}}s_{p}(t).$ (2)
V-A-4) Eavesdropper Setup
We consider two different knowledge levels for Eve:
  1. 1.
    Full knowledge, where Eve shares the identical knowledge with Bob about the system design, including the VENENA framework, the protocol, the data format, the parameter specifications, the image classification network used by Bob, as well as the mapping table; and
  2. 2.
    partia knowledge, where Eve knows the basic visual encryption scheme and the mapping table, but is not aware of the VENENA framework.
In this article, we focus on the case where Eve to be consistently eavesdropping. For readers who are interested in sophisticated dynamic eavesdropping strategies, we refer to our related work [5].
V-A-5) Radio Setup
We set the channel bandwidth to $1\text{\,}\mathrm{MHz}$ and the gross transmission bit rate to $1\text{\,}\mathrm{Mbit}$$\mathrm{s}$. To reflect realistic multipath wireless scenarios, we consider Rayleigh fading non-line-of-sight (NLOS) channels for both Bob and Eve, with a higher mean path loss in the eavesdropper channel than that in the legitimate channel. The power density of additive white Gaussian noise (AWGN) is $174\text{\,}\mathrm{dBW}$$\mathrm{Hz}$ at both receivers.
V-B VALIDATION TEST
Table 4: Benchmark results: message perception accuracy
VENENA, 75% mixing NVE, full power NVE, 75% power NVE, conventional PLS*
Bob 93.43% 97.22% 96.91% 93.42%
Eve (full knowledge) 51.12% 88.96% 86.82% 60.22%
Eve (partial knowledge) 5.18% 89.14% 86.66% 60.22%
Figure 2: Sensitivity of message perception accuracy versus mixing ratio $\alpha$ under different transmission power and eavesdropping channel conditions.
First we conducted a validation test to demonstrate the capability of the VENENA framework to securely transmit confidential messages. We considered the transmission power budget $P_{\Sigma}=$100\text{\,}\mathrm{mW}$\$, mean legitimate channel gain $\overline{z}_{\mathrm{Bob}}=$-85\text{\,}\mathrm{dB}$\$, mean eavesdropping channel gain $\overline{z}_{\mathrm{Eve}}=$-95\text{\,}\mathrm{dB}$\$, and the following four schemes for benchmarking:
  1. 1.
    VENENA, 75% mixing: $\alpha=0.75$, $P_{\Sigma}=$100\text{\,}\mathrm{mW}$\$, which provides Bob a small SINR of $3\text{\,}\mathrm{dB}$ for the poison mask decoding, corresponding to channel capacity of $1.58\text{\,}\mathrm{MHz}\text{/}\mathrm{s}$ that slightly exceeds the transmission rate;
  2. 2.
    Naive Visual Encryption (NVE), full power: $\alpha=1$, $P_{\Sigma}=$100\text{\,}\mathrm{mW}$\$;
  3. 3.
    NVE, 75% power: $\alpha=1$, $P_{\Sigma}=$75\text{\,}\mathrm{mW}$\$; and
  4. 4.
    NVE, conventional PLS baseline: $\alpha=1$, transmission power adjusted to the level that Bob achieves the same perception accuracy as in the VENENA-75%-mixing scheme.
We carried out the Monte-Carlo test with 1000010~000 independent trails under each scheme, where Bob, full-knowledge Eve, and partial-knowledge Eve are measured regarding their and capability of accurately perceiving the confidential message. The results are summarized in Tab. 4. It can be observed that compared to all NVE baselines including that assisted by conventional PLS, the VENENA framework can effectively deceive Eve even when it has full knowledge about the system design, while maintaining a high message perception accuracy for Bob.
These optimization objectives can be realized through dynamic mixing strategies in practical deployment. The optimal mixing ratio $\alpha$ depends on both legitimate and eavesdropping channel conditions. While Bob’s channel condition can be directly measured through standard channel estimation techniques, Eve’s channel statistics can be estimated based on geographical security zones, historical eavesdropping patterns, or worst-case assumptions for conservative design. A practical implementation would pre-compute performance lookup tables through large-scale simulation campaigns covering diverse channel condition combinations. During operation, the mixing ratio is adaptively selected to maximize the chosen objective (e.g., accuracy gap or Eve’s failure rate) while maintaining Bob’s minimum accuracy requirement. This channel-sensing-based dynamic mixing enables VENENA to adapt to time-varying wireless environments.
V-C SENSITIVITY TEST
We further conducted a sensitivity test to evaluate the sensitivity of the performance of VENENA against the transmission power and the eavesdropping channel condition under different mixing ratios. Similarly to the verification test, we considered Bob with mean channel gain $\overline{z}_{\mathrm{Bob}}=$-85\text{\,}\mathrm{dB}$\$, Eve with full knowledge, and Eve with partial knowledge, and investigate their performance in three different scenes:
  1. 1.
    Scene 1: $P_{\Sigma}=$100\text{\,}\mathrm{mW}$\$, $\overline{z}_{\mathrm{Eve}}=$-95\text{\,}\mathrm{dB}$\$;
  2. 2.
    Scene 2: $P_{\Sigma}=$200\text{\,}\mathrm{mW}$\$, $\overline{z}_{\mathrm{Eve}}=$-95\text{\,}\mathrm{dB}$\$; and
  3. 3.
    Scene 3: $P_{\Sigma}=$100\text{\,}\mathrm{mW}$\$, $\overline{z}_{\mathrm{Eve}}=$-90\text{\,}\mathrm{dB}$\$.
In all three scenes we measured the message perception accuracy of Bob and Eve under different mixing ratios ranging from $0.5$ to $0.95$. For each scene and each mixing ratio, we conducted $1000$ independent trails.
The results are depicted in Fig. 2, from which we can observe that Eve’s performance monotonically decreases with the raising mixing ratio $\alpha$ in all three scenes, as long as it is not aware of the VENENA framework. For both Bob and Eve with full knowledge, the performance is generally concave of the power mixing ratio in all three scenes, with the maximum perception accuracy achieved around $75\%$ mixing. The optimal specification of the overall system, however, depends on the optimaziation objective, e.g. the perception accuracy gap between Bob and Eve, or the perception failure rate of Eve under the constraint of Bob’s mimimum perception accuracy.
For a more intuitive illustration of VENENA’s descriminative deception on Eve, we present in Fig. 3 the receiver operating characteristic (ROC) curves showing the deception rate of Bob and full-knowledged Eve upon different mixing ratios in the three scenes.
Figure 3: Deception rates of Bob and full-knowledge Eve, respectively, under different mixing ratios in three scenes.
6 DISCUSSION
Towards a more practical deployment of the VENENA framework in realistic communication scenarios, some discussions regarding the efficiency aspects are necessary.
VI-A ENCRYPTION CODING EFFICIENCY
The encryption coding efficiency of the VENENA framework is determined by the size of the image and the amount of available semantic tags. For the demonstrational setup we have used in this article, the coding rate is as low as $\log_{2}^{10}/(32\times 32\times 3\times 8)\approx 1.35\times 10^{-4}$, which is applicable only for low-speed communication with high secrecy requirements. To raise the efficiency for high-speed data transmission scenarios, several approaches can be considered:
  1. 1.
    Smaller images: Reducing image dimensions directly decreases transmission overhead, though this will degrade the decoding accuracy of the classification network.
  2. 2.
    More semantic tags: Introducing additional semantic tags to the mapping table increases coding rate logarithmically, but requires a larger image database and more poisoning networks to be trained.
  3. 3.
    Multi-tag images: Using images containing multiple semantic tags simultaneously (e.g., an image showing both an airplane and a bird) significantly increases information density. However, this substantially complicates both database construction–requiring systematic labeling of multi-tag combinations–and poisoning network training, as the network must achieve targeted misclassification across multiple semantic dimensions simultaneously.
  4. 4.
    Multi-modal encoding: Extending beyond visual content to incorporate audio or text modalities enables parallel transmission of multiple semantic channels, multiplicatively increasing the effective coding rate. This requires developing poisoning networks for each modality and ensuring cross-modal consistency.
  5. 5.
    Semantic-preserving compression: Pre-processing images with neural compression techniques can reduce transmission overhead significantly while maintaining classification accuracy, achieving superior rate-distortion tradeoffs compared to traditional codecs without requiring additional database or network resources. It shall be noted that this approach introduces additional computational complexity and delay at both the transmitter and receiver.
VI-B COMPLEXITY AND SCALABILITY
By increasing the number of semantic tags, the encryption coding efficiency can be linearly improved. However, this will also raise the space complexity, as it also requires linear increases in both the size of image database and the set of poisoning networks (since each network is specified to a dedicated semantic pair). Although the complexities of runtime encoding, poisoning, transmission, and decoding are not impacted by the semantic tag number $n$, a $\mathcal{O}(n)$ space complexity in database, a $\mathcal{O}(n^{2})$ space complexity in poisoning networks, and a $\mathcal{O}(n^{2})$ time complexity in network training are expected. Advanced techniques such as data distillation [23] and network pruning [24] shall be considered, in this context, to improve the scalability of the VENENA framework for practical deployment with a large number of semantic tags.
VI-C CHANNEL CONDITION ASSUMPTIONS
VENENA’s security guarantee relies on the assumption that Eve experiences statistically inferior channel conditions compared to Bob. While this assumption does not hold universally, it reflects realistic deployment scenarios in modern wireless networks [3]. Fixed infrastructure transmitters (e.g., base stations) typically operate within secrecy-protected zones that prevent eavesdroppers from maintaining close proximity, either through physical security measures or detection mechanisms [20, 21]. For mobile receivers, anonymity provisions in contemporary wireless protocols make precise localization challenging for adversaries, hindering their ability to consistently position themselves advantageously [22]. Furthermore, directional beamforming towards legitimate receivers naturally degrades eavesdropper channel quality when they cannot collocate with intended recipients. These factors collectively justify the channel asymmetry assumption for many practical scenarios including cellular downlink, wireless backhaul, and device-to-device communications. Nevertheless, scenarios where adversaries achieve sustained proximity to either transmitter or receiver (e.g., compromised infrastructure, targeted surveillance with dedicated resources) represent limitations of the current framework and motivate future research on adaptive security mechanisms that degrade gracefully under adverse channel conditions.
7 CONCLUSION AND OUTLOOKS
In this work, we have proposed VENENA, a novel framework that combines physical layer deception with visual encryption and image poisoning for secure wireless transmission of semantics. By integrating vision transformer-based poisoning networks with power-domain multiplexing, we achieve effective deception of eavesdroppers while maintaining reliable communication with legitimate receivers.
Experimental validation demonstrates that VENENA maintains above 93% message perception accuracy for legitimate receivers while reducing eavesdroppers’ success rate by more than 40%–even when adversaries possess full system knowledge. This quantifiable benchmark confirms the framework’s active defense capability and practical robustness for 6G semantic communication scenarios.
The sensitivity analysis further identifies optimal system configurations, particularly regarding the power mixing ratio , and reveals VENENA’s adaptability to diverse channel conditions. The discussion on efficiency, scalability, and channel assumptions provides guidance for future extensions, including multi-modal semantic encryption, adaptive mixing, and cross-layer deception mechanisms.
With new possibilities in wireless physical layer security opened in this work, variate potential future topics are suggested towards future study. These are including but not limited to:
  1. 1.
    the extension of the proposed framework to more complex visual content, such as high-resolution images that contain multiple semantic tags in each;
  2. 2.
    the adoption to flexible semantic encoding that uses other forms of semantic information carriers, such as text or audio; and
  3. 3.
    the development of dynamic adaptation strategies for challenging channel conditions.